OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−2, −2) and (2, 4)?

OpenStudy (3psilon):

Use the slope formula \[\frac{ y_{2}-y_{1} }{ x_{2}-x_{1} }\]

OpenStudy (anonymous):

Oh do this again? Okay got it hahah

OpenStudy (3psilon):

\[\frac{ 4-(-2) }{2-(-2) }\]

OpenStudy (3psilon):

Yes . Get it ? :)

OpenStudy (anonymous):

I do! thanks!

OpenStudy (3psilon):

But wait that's the slope. You need the y intercept which is the b in y = mx +b

OpenStudy (anonymous):

oh.

OpenStudy (3psilon):

So we found the slope m . Now just plug in One of the two coordinates and the slope into y = mx + b and solve for b. Let me know if you need more help

OpenStudy (anonymous):

I need more help forsure

OpenStudy (3psilon):

Ok did you find the slope m?

OpenStudy (anonymous):

finding it now.

OpenStudy (3psilon):

Ok let me know what you get

OpenStudy (anonymous):

I got 3/2?

OpenStudy (3psilon):

Good job! Now we use the slope intercept formula y = mx + b

OpenStudy (3psilon):

pick either (−2, −2) or (2, 4) to plug in for x and y

OpenStudy (3psilon):

And you'll be able to solve for the y - intercept b

OpenStudy (anonymous):

so -2=3/2(-2)+b? lo

OpenStudy (3psilon):

Perfect! Now just solve for b and you have your answer!

OpenStudy (anonymous):

how though?!

OpenStudy (anonymous):

I just cant find an answer...

OpenStudy (3psilon):

Just isolate the term b

OpenStudy (3psilon):

b is the y intercept

OpenStudy (anonymous):

Is -1=3/2(-2) it?

OpenStudy (3psilon):

nope I'll draw it out for you. one sec

OpenStudy (anonymous):

ok

OpenStudy (3psilon):